OpenStudy (gamer56):

what is the solution to the equation? 3x -6= -2x + 9 A.3/5 B.15 C.3 D.5/2

OpenStudy (gamer56):

@KlOwNlOvE

OpenStudy (anonymous):

i got this ready?

OpenStudy (gamer56):

?

OpenStudy (anonymous):

C. 3

OpenStudy (anonymous):

C!

OpenStudy (anonymous):

dude

OpenStudy (anonymous):

ur a human calc

OpenStudy (gabylovesyou):

http://www.wolframalpha.com/input/?i=3x+-6%3D+-2x+%2B+9 so the answer is C :)

OpenStudy (gamer56):

@cambrige sometimes a human calculator needs help also.

OpenStudy (anonymous):

seriously??

OpenStudy (anonymous):

sometimes i need help.... All depends on the question

OpenStudy (anonymous):

I assumed uur like a wizard at maths

OpenStudy (like1234):

Simplifying 3x + -6 = -2x + 9 Reorder the terms: -6 + 3x = -2x + 9 Reorder the terms: -6 + 3x = 9 + -2x Solving -6 + 3x = 9 + -2x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '2x' to each side of the equation. -6 + 3x + 2x = 9 + -2x + 2x Combine like terms: 3x + 2x = 5x -6 + 5x = 9 + -2x + 2x Combine like terms: -2x + 2x = 0 -6 + 5x = 9 + 0 -6 + 5x = 9 Add '6' to each side of the equation. -6 + 6 + 5x = 9 + 6 Combine like terms: -6 + 6 = 0 0 + 5x = 9 + 6 5x = 9 + 6 Combine like terms: 9 + 6 = 15 5x = 15 Divide each side by '5'. x = 3 Simplifying x = 3 http://www.geteasysolution.com/3x-6=-2x+9

OpenStudy (anonymous):

3x-6=-2x+9 5x=15 (got x's on one side values on the other) x=3

OpenStudy (like1234):

the answer is C

OpenStudy (gamer56):

Thanks guys :D

OpenStudy (anonymous):

that's a biga** simplification @like1234